OpenStudy (domcul):

What is the slope-intercept form of the equation of a line that passes through (5, -4) and has a slope of 3/4?

OpenStudy (luckylu):

y= mx+b is the formula

OpenStudy (domcul):

okay

OpenStudy (luckylu):

since 3/4 is the slope that is m

OpenStudy (domcul):

okay so it would be y=3/4x + 1/4?

OpenStudy (luckylu):

where did 1/4 come from?

OpenStudy (domcul):

i dont really know its just one of the choices

OpenStudy (luckylu):

can I see the choices

OpenStudy (domcul):

A: y=3/4x+1/4 B: y=3/4x+31/4

OpenStudy (luckylu):

yeah go with a

OpenStudy (domcul):

Thank you!